OpenStudy (anonymous):

simplify the expression, cos(90° - θ) I am not sure what to do

OpenStudy (campbell_st):

 use trig ratios looking at the diagram \[\cos(90 - \theta) = \frac{b}{c}\] this ratio is the same as \[\sin(\theta) = \frac{b}{c}\]